OpenStudy (anonymous):

How to solve sin 2λ ?

OpenStudy (unklerhaukus):

that is not an equation

OpenStudy (anonymous):

sin 2λ = 0 2λ = n.pi λ = n.pi / 2 ----->>>> How do I get this...???

OpenStudy (unklerhaukus):

\[sin (2\lambda) =0\] \[2\lambda = arcsin(0)\] \[ \lambda = { arcsin(0) \over 2} \]

OpenStudy (unklerhaukus):

\[arcsin (0) =0+n\pi\] where n is an integer

OpenStudy (unklerhaukus):

graph y = arcsin(x)

OpenStudy (anonymous):

How about arccos(0) ?

OpenStudy (anonymous):

\[arccos{0}=\frac{\pi}{2}+n\pi\] where n is an integer